OpenStudy (anonymous):

Evaluate __________. A. 9 B. 362,880 C. 126 D. 3,024

OpenStudy (anonymous):



OpenStudy (anonymous):

\[_9P_4\]?

OpenStudy (anonymous):

Yes!!!

OpenStudy (anonymous):

count out four places and multiply \[9\times 8\times 7\times 6\]

OpenStudy (unklerhaukus):

\[_nP_k= \frac{n!}{(n-p)!}\]

OpenStudy (anonymous):

What do u mean count out 4 places

OpenStudy (anonymous):

i mean start with \(9\)

OpenStudy (anonymous):

3,024?

OpenStudy (anonymous):

then times 8, times 7, times 6 use 4 numbers

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

\[_9P_4= \frac{9!}{(9-4)!}\]\[=\frac{9!}{5!}\]\[=\frac{9\times8\times7\times6\times5!}{5!}\]\[=9\times8\times7\times6\]

OpenStudy (anonymous):

@UnkleRhaukus wrote the formula, but imho you should not use it as written the number after the P tells you how many numbers you are going to multiply together

OpenStudy (anonymous):

so for example \[_{10}P_3=10\times 9\times 8\]

OpenStudy (anonymous):

Okay. thank you.